Please help me:
2. If cos x=1/2, then sin x= ?
3. If cotx=1, then x= ?
Thanks

Answer by stanbon(57228)

2. If cos x=1/2, then sin x= ?
----
If cos(x) = 1/2 then x=1 and r = 2
Then y = sqrt(r^2-x^2) = sqrt(4-1) = sqrt(3)
----
But y can be positive or negative.
----
So sin(x) = y/r = sqrt(3)/2 or sin(x) = y/r = -sqrt(3)/2
=============================================================
3. If cotx=1, then x= ?
---
x = 45 degrees or x = 180+45 = 225 degrees
================================================
Cheers,
Stan H.